OpenStudy (anonymous):

Logarithm Help needed, writing question below :) Just need someone to double-check an answer

OpenStudy (anonymous):



OpenStudy (anonymous):

@KevinOrr

OpenStudy (phi):

no. you can use a calculator to show it does not work

OpenStudy (anonymous):

K, the other answer choices r:

OpenStudy (anonymous):

I think I'm doing a mistake*

OpenStudy (phi):

take the log (base 10 or base e ) of both sides log ( 3^(1-2x)) = log(2) (1-2x) log(3) = log(2) 1-2x = log(2)/log(3) 1- log(2)/log(3) = 2x (1-log(2)/log(3)) /2 = x

OpenStudy (jdoe0001):

$$\bf 3^{1-2x} = 2\\ log_3(3^{1-2x}) = log_32\\ 1-2x = \cfrac{log(2)}{log(3)}\\ \cfrac{ 1+\frac{log(2)}{log(3)}}{ 2} = x $$

OpenStudy (phi):

I skipped some steps on the algebra... I assume you follow ?

OpenStudy (anonymous):

No wonder I did [1- log(2)/log(3)]/2 for some reason Thx :) gonna figure it out now

OpenStudy (jdoe0001):

hmmm, wait a minute, you're right... should be negative :(

OpenStudy (anonymous):

and i flipped the logs around,

OpenStudy (jdoe0001):

$$\bf 3^{1-2x} = 2\\ log_3(3^{1-2x}) = log_32\\ 1-2x = \cfrac{log(2)}{log(3)}\\ \cfrac{ 1-\frac{log(2)}{log(3)}}{ 2} = x $$

OpenStudy (phi):

that looks good. [1- log(2)/log(3)]/2 is your calculator broken?

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

lol it's hard when u have 2 use the internet as ur calculator and i got x~.185

OpenStudy (phi):

yes. 0.185

OpenStudy (anonymous):

Thx a ton to both of u! :)